OpenStudy (anonymous):

how to solve quadratic equation? 1/10+1/15=1/x How I solved: (3/3)1/10 + (2/2)1/15 = 1/x (30/30) = 3/30+2/30=5/30 x = 30 other way of solving: 1/10 + 1/15 = 1/x Divide each denominator by 5: 1/2 + 1/3 = 1/5 x = 5 Is either of these solutions correct, or is there another way?

OpenStudy (anonymous):

as LCM of 10 and 15 is 30 so, multiply both sides by 30 first..

OpenStudy (anonymous):

not a quadratic equation, and your steps are odd you can not remove the x from the denominator, unless you multiply by x to the entire equation you cannot add the denominators together

OpenStudy (anonymous):

@fiveatheart59 got it now..??

OpenStudy (anonymous):

@completeidiot agree with you.. he did it in a strange way..

OpenStudy (anonymous):

@Kashan yes, thank you

hero (hero):

\[\frac{1}{10} + \frac{1}{15} = \frac{1}{x}\] Multiply 1/10 by 3/3 to get 3/30 Multiply 1/15 by 2/2 to get 2/30: \[\frac{3}{30} + \frac{2}{30} = \frac{1}{x}\] Combine fractions on the left side to get: \[\frac{5}{30} = \frac{1}{x}\] Reduce 5/30 to 1/6: \[\frac{1}{6} = \frac{1}{x}\] Therefore, x = 6